# Learning Optimal Control with Stochastic Models of Hamiltonian Dynamics

### v2

@@ Abstract
Optimal control problems can be solved by applying the Pontryagin maximum principle and then solving for a Hamiltonian dynamical system. In this paper, we propose novel learning frameworks to tackle optimal control problems. By applying the Pontryagin maximum principle to the original optimal control problem, the learning focus shifts to reduced Hamiltonian dynamics and corresponding adjoint variables. The reduced Hamiltonian networks can be learned by going backward in time and then minimizing loss function deduced from the Pontryagin maximum principle’s conditions. The learning process is further improved by progressively learning a posterior distribution of reduced Hamiltonians, utilizing a variational autoencoder which leads to more effective path exploration process. We apply our learning frameworks to control tasks and obtain competitive results.
Keywords: Reinforcement learning, Pontryagin maximum principle, Hamiltonian neural network
@@ 1 Introduction
Learning optimal control solution models for unknown dynamical systems, where an objective cost functional is optimized over space-time state-action sequences, is known to be computationally similar to convergently learning an optimal state-action policy in reinforcement learning, actioning on optimal value paths with maximum reward or minimum regret (Todorov (2006)). The solutions to such reinforcement learning methods further involve stable numerical solution techniques for high-dimensional and high variance constrained optimization. Several reinforcement learning algorithms have been devised to attempt to compute the optimal control paths through guided search and policy learning (Abbeel et al. (2006); Deisenroth and Rasmussen (2011); Levine and Abbeel (2014); Heess et al. (2015); Mnih et al. (2015)).
In this paper, we take a more comprehensive approach. We rely on both the physical nature of the unknown dynamical system and the necessary requirements of the optimal control sequences or trajectories subjected to these physical dynamics (Boltyanskiy et al. (1962)). We learn the underlying reduced Hamiltonian dynamics and then learn to utilize this data dependent reduced Hamiltonians to optimally control the dynamical system for targeted objectives. We focus on the discrete-time with uneven time steps and continuous-time versions of the optimal control optimization problem. By using the Pontryagin maximum principle, we reduce the optimal control problem to the data specific form of the Hamiltonian dynamics learning. The control is now encoded into the adjoint variable, and the new reinforcement learning task is to learn through this adjoint variable and the reduced Hamiltonian. This is fundamentally different from optimal policy driven reinforcement learning, which focuses on deriving optimal action and value functions (Mnih et al. (2015); Gu et al. (2016)). In particular, we first learn a correct reduced Hamiltonian of the controlled dynamics, obtained by the application of the Pontryagin maximum principle to the original optimal control problem. The reduced Hamiltonian deep network can be trained by going backwards in time and from minimizing a loss function deduced from the Pontryagin maximum principle’s conditions. Moreover, the learning process is improved through variational inference and use of a variational autoencoder by progressively learning a posterior distribution of reduced Hamiltonians, which leads to a filtration of the generalized coordinates during the neural network training.
@@ 1.1 Related works
Model-based reinforcement learning algorithms. Several model-based algorithms (Wang et al. (2019)) have been developed for reinforcement learning. Unlike its model-free counterparts (Mnih et al. (2016, 2015); Schulman et al. (2015, 2017)), model-based approachs attempt to model the outside environment and then use suitable mathematical formulation to facilitate the finding of agent’s optimal actions. Two popular model-based approaches are the Dyna-style algorithms (Sutton (1990)) that model the environment to generate imaginary data for actual policy training, and policy search with backpropagation through time. The latter approach with notable algorithms such as PILCO by Deisenroth and Rasmussen (2011), computes the analytic gradient (closed-form formulas) of the reinforcement learning objective with respect to the policy via explicit modeling of the problem/environment.
PMP-based framework on optimal control. Many works are developed based on the Pontryagin maximum principle (PMP) to tackle optimal control problems. For instance, Pontryagin Neural Networks (PoNNs) by D’Ambrosio et al. (2021) use PMP to transform the original problem into a Two-Point Boundary Value ordinary differential equation (ODE). PoNNs then use neural-network regression to learn such an ODE. Pontryagin Differentiable Programming by Jin et al. (2020) and AI Pontryagin by Böttcher et al. (2022), on the other hand, use PMP to derive the exact analytic gradient of a certain loss function with respect to the parameters of the control to be learned. This method is similar to the second popular approach of the model-based reinforcement learning. Notice that all of these works use PMP only to derive a closed form formula that are then used to train the actual learning architecture. In these cases, PMP is not actively incorporated in the learning framework itself.
PMP-based deep learning framework. There is also a class of research works (Li et al. (2017b); Zhang et al. (2019)) that use iterative PMP-based algorithms to train the Hamiltonian and then derive the optimal control variable. However, this line of work aims to improve the traditional supervised deep learning training rather than focusing on the optimal control tasks themselves. The goal is also on optimizing the parameters of deep learning architectures rather than the optimal paths of the control problem.
Physics-informed dynamic learning.  Several physics-informed learning architectures aim at learning hidden dynamics by incorporating physics biases. Deep Lagrangian Networks (DeLaN) by Lutter et al. (2019), for example, tries to inject Lagrangian mechanics bias into the supervised learning process to improve the performance on difficult trajectories. Sympletic-ODE Net by Zhong et al. (2019), on the other hand, learns the unknown dynamics of a controlled system with Hamiltonian bias. Similar to our framework, Sympletic-ODE also exploits the Hamiltonian dynamics for more accurate learning:
@@
Neural ODE.  Similar to our work, Sympletic-ODE Net builds on the Neural ODE architecture by Chen et al. (2018), which proposes to model a neural network by an ordinary differential equation and learns the rate of change $\frac{dy}{dx}=f(x,y,\theta)$ rather than the function $f$ itself. The backpropagation of NeuralODE is based on the adjoint method with a backward ordinary differential equation on the adjoint states $a(t)=\frac{dL}{dh(t)}$. Comparing with traditional DNNs, Neural ODE (Dupont et al. (2019)) allows continuous time-series modeling and doesn’t require discretizing observation and emission intervals.
@@ 1.2 Contributions
We propose novel learning-based frameworks to solve general optimal control problems. Our framework actively incorporates Pontryagin maximum principle (PMP) into the learning process. Our contributions include:
@@
    A two phases-learning framework called NeuralPMP (see section 4) that actively uses Pontryagin maximum principle in the training loss function and the learning process. The two phases utilize a variational autoencoder to prevent a static learning while helping the model learn the correct Hamiltonian and the corresponding dynamics.
@@ 1.3 Organization
The rest of the paper is organized as follows. Section $2$ briefly introduces the mathematical background including optimal control formulation, the Pontryagin maximum principle (PMP), and the definitions of reduced Hamiltonian. Section $3$ proposes an algorithm that incorporates PMP into the training loss to solve discrete-time control problems. Sections 4 builds on the limitation of the approach in Section $3$ and present in details a two phase-learning framework called NeuralPMP for continuous-time optimal control. Section $5$ provides experimental results of NeuralPMP on classical control tasks. The final section $6$ concludes our work and outlines future research directions.
@@ 2.1 Optimal control formulation
For simplicity, we work on Euclidean spaces. Our concept can be generalized to a general Banach space. Assume we have state space $X$, a bounded subset of $\mathbb{R}^{m}$, with dual $X^{*}$, which can be identified with $\mathbb{R}^{m}$, and a control space $U$, which is another bounded subset of $\mathbb{R}^{n}$. Here $m\in\mathbb{N}$ and $n\in\mathbb{N}$ are the state space and control space dimensions. Given a terminal time $T$, we need to find an optimal control path (also called policy) $u^{*}$ so that $u^{*}(t)\in U$ for all time $t\in[0,T]$ and such $u^{*}$ minimizes the following cost functional:
@@ 2.3 Pontryagin maximum principle
The Pontryagin maximum principle (Kirk (1971)) provides the necessary condition for an optimal solution of the general optimal control problem.
Theorem 1
@@ 3 Discrete-time Hamiltonian dynamics learning
Before jumping into our main framework in section 4, we propose a discrete-time framework to learn the discretized reduced Hamiltonian dynamics, and analyze its advantage over the popular policy-based methods on a specific, irregular linear quadratic control problem.
@@ 3.1 Discrete framework with PMP-based loss
Problem. Suppose an agent $\mathcal{A}$, initially starts at $q_{0}$, needs to perform a sequence of actions $u_{0},u_{1},\cdots,u_{T-1}$ to perform at times $t_{0}=0,t_{1},\cdots,t_{T-1}$ in order to maximize the cost functional $J(q,u)$ introduced in eq. 6.
Solution. From the theory in section 2, we aim to build a parameterized neural network $F_{\theta}$ with parameter $\theta$ that maps from initial state $q_{0}$ to the sequence of optimal actions $(u_{0},u_{1},u_{2},\cdots,u_{T-1})$ by minimizing a Pontryagin maximum principle-based loss function. The training includes 3 steps:
@@
    The adjoint variables $p_{k}$ are then calculated backward in time $p_{k-1}=p_{k}+\Delta_{k}H_{q}(q_{k},u_{k},p_{k})$, where $H$ is the Hamiltonian defined in eq. 8.
@@
The dynamical system equations eq. 10 and eq. 11 are effectively discretized in the last two steps of the forward pass. We aim to train a network that preserves the last condition eq. 12 of Pontryagin maximum principle (PMP): $\partial_{u}H(x(t),p(t),u(t))=0$. As a result, the loss function for the (policy) network $F_{\theta}$ is defined as:
$L(\theta)=\sum_{k=0}^{T}H_{u}(q_{k},u_{k},p_{k})$ (16)
@@ 3.2 Special linear quadratic control problem
subject to the linear dynamics $\dot{a}=Aa+Bu$. For the special case where state space $X=\mathbb{R},A=0,B=1,R=1,Q=0$, the LQR problem is equivalent to the task when a learning agent needs to control its velocity to get as close to the origin as possible even though its initial position can be far from 00. We train the policy $F_{\theta}$ based on the PMP loss on $100$ iterations with 1600 random samples per iterations. The discrete-time horizon is chosen to be $[0,1,2,4,8]$. Notice that the times are not equally spaced, and the values of timestamps are taken into the account in this model. The plot (see fig. 1a) show that the training process converges after $70$th iteration.
With the actions produced from trained policy $F_{\theta}$, the trained car agent quickly decreases the velocity in order to get to 00 as quick as possible at time 00. At the subsequent times, the car learns to stay where it was. This is the optimal behavior given that the car needs to minimize its cumulative distance to 00 over the time period from 00 to $8$ despite its starting position (see fig. 1b).
Policy-based algorithms in reinforcement learning learn a policy function $\pi$ mapping from $q_{t}$ to the optimal action $u_{t}$. In this special case, we show that no such reasonable function $\pi$ exists. Assume by contradiction, we have such a function $\pi$ so that $u_{t}=\pi(q_{t})$ for each $t$. For $t\geq 1$, $\pi(q_{t})=u_{t}$ must be 00 because once the car get to the origin, it should stay there forever. It implies that the policy function is approximately zero function. However, the car must increase or decrease its speed to get to 00 as quickly as possible right after it starts. Thus, for a starting value $q_{0}\neq 0$, $u_{0}$, which is approximated by $\pi(q_{0})=0$, needs to be non-zero. Contradiction.
@@ 4 Continuous Hamiltonian dynamics learning with forward-backward variational autoencoder
The previous training framework in section 3 allows us to gain important ideas in actively incorporating PMP principles into the deep learning framework. Nonetheless, the mapping $F_{\theta}$ introduced earlier can be very costly for the time horizon with many steps $t_{1},\cdots,t_{T}$ for large $T$, as it maps to the entire sequence of actions. Moreover, even though the framework can handle uneven time steps, it still requires the discretization of the underlying dynamics, and may not be applicable to a general continuous-time controlled dynamical system. In this section, by also incorporating PMP conditions into the training process, we build a learning framework for continuous-time optimal control problem with a focus on learning the reduced Hamiltonian function eq. 13 and the corresponding Hamiltonian flow.
@@ 4.1 Reduced Hamiltonian learning with PMP-based training loss
    A $P$-net $P_{\phi}$ that takes input as the starting state $q_{0}$ and outputs the adjoint variable $p_{0}$. The $p_{0}$ is necessary as the Hamiltonian $h_{\theta}$ also needs an adjoint variable as input in addition to the state variable.
@@
[…]
@@
    Use NeuralODE (Chen et al. (2018)) to sample the generalized coordinates $(q_{t},p_{t})_{t\geq 0}$ from the Hamiltonian flow defined by equations eq. 14 and eq. 15 with $h=h_{\theta}$. Then for each generalized coordinates $(q_{t},p_{t})$ on the Hamiltonian flow, the 5-tuple $(q_{t},p_{t},\hat{u}(q_{t},p_{t}),\hat{f}(q_{t},\hat{u}_{t}),l(q_{t},\hat{u}_{t}))$ is added to $\mathcal{M}$.
@@
    where $p_{T}$ is the terminal adjoint variable after running $(q,P_{\phi}(q))$ through Hamiltonian dynamics specified by $h_{\theta}$. The variables $\alpha_{1},\alpha_{2}$ and $\beta_{1}$ are hyper-parameters and are optimizable by several techniques: Bayesian optimization (Snoek et al. (2012)), Tree-Structured Parzen estimator (Bergstra et al. (2011)), or bandit algorithms (Li et al. (2017a)). For efficiency and simplicity, we currently employ grid search with early termination. The first two terms are based on the PMP condition eq. 9, while the last term is used to learn the reduced Hamiltonian correctly.

### v3

@@ Abstract
In this paper, we propose novel learning frameworks to tackle optimal control problems by applying the Pontryagin maximum principle and then solving for a Hamiltonian dynamical system. Applying the Pontryagin maximum principle to the original optimal control problem shifts the learning focus to reduced Hamiltonian dynamics and corresponding adjoint variables. Then, the reduced Hamiltonian networks can be learned by going backwards in time and then minimizing loss function deduced from the Pontryagin maximum principle’s conditions. The learning process is further improved by progressively learning a posterior distribution of the reduced Hamiltonians. This is achieved through utilizing a variational autoencoder which leads to more effective path exploration process. We apply our learning frameworks called NeuralPMP to various control tasks and obtain competitive results.
keywords:
Reinforcement learning, Pontryagin maximum principle, Hamiltonian neural network
@@ 1 Introduction
Learning optimal control solution models for unknown dynamical systems, where an objective cost functional is optimized over space-time state-action sequences, is known to be computationally similar to convergently learning an optimal state-action policy in reinforcement learning. The actions on optimal value paths achieve maximum reward or minimum regret [26]. The solutions to such reinforcement learning methods involve stable numerical solution techniques for high-dimensional and high variance constrained optimization. Several reinforcement learning algorithms have been devised attempting to compute the optimal control paths through guided search and policy learning [1, 7, 15, 11, 20, 2].
In this paper, we take a more comprehensive approach. We rely mainly on two things: the physical nature or environment of the unknown dynamical system and the necessary requirements of the optimal control sequences or trajectories subjected to these physical dynamics [21]. We learn the underlying reduced Hamiltonian dynamics and then utilize this data-dependent reduced Hamiltonians to optimally control the dynamical system for targeted objectives. We focus on the discrete-time with uneven time steps and continuous-time versions of the optimal control optimization problem. Through Pontryagin maximum principle, we reduce the optimal control problem to the data specific form of the Hamiltonian dynamics learning. The control is now encoded in the adjoint variable. Therefore, the new reinforcement learning task is to learn through this adjoint variable and the reduced Hamiltonian. This is fundamentally different from the optimal policy driven reinforcement learning, which focuses on deriving optimal action and value functions [20, 10]. In particular, we first learn a correct reduced Hamiltonian of the controlled dynamics. This is obtained via the application of the Pontryagin maximum principle to the original optimal control problem. The reduced Hamiltonian deep network can be trained by going backwards in time and minimizing a Pontryagin maximum principle-based loss function. Moreover, the learning process is improved through variational inference and by learning a posterior distribution of the reduced Hamiltonians. This process leads to a filtration of the generalized coordinates during the neural network training.
@@ 1.1 Related works
Model-based reinforcement learning algorithms. Several model-based algorithms [27] have been developed for reinforcement learning. Unlike their model-free counterparts [19, 20, 22, 23], model-based approaches model the outside environment and then use suitable mathematical formulations to facilitate the finding of agent’s optimal actions. Two popular model-based approaches include the Dyna-style algorithms [25] and algorithms focusing on policy search with backpropagation through time. While the first approach (Dyna-style algorithms) models the environment to generate imaginary data for actual policy training, the second approach, with notable methods such as PILCO [7], computes the analytic gradient (closed-form formulas) of the reinforcement learning objective with respect to the policy via explicit modeling of the problem/environment.
PMP-based framework on optimal control. Many works are developed based on the Pontryagin maximum principle (PMP) to tackle optimal control problems. For instance, Pontryagin Neural Networks (PoNNs) [9] use PMP to transform the original problem into a Two-Point Boundary Value ordinary differential equation (ODE). PoNNs then use neural-network regression to learn such an ODE. On the other hand, Pontryagin Differentiable Programming [12] and AI Pontryagin [4] use PMP to derive the exact analytic gradient of a certain loss function with respect to the parameters of the control to be learned. This method is similar to the second popular approach of the model-based reinforcement learning. We remark that all of these works leverage PMP only to derive closed-form formulas that can then be used to train the actual learning architecture. In these cases, PMP is not actively incorporated in the learning framework.
PMP-based deep learning framework. There is also a line of research works [17, 28] that use iterative PMP-based algorithms to train the Hamiltonian and then derive the optimal control variable. However, this line of works aims to improve the traditional supervised deep learning training rather than solving the optimal control tasks themselves. Their goals also focus on optimizing the parameters of deep learning architectures rather than the optimal paths of the control problem.
Physics-informed dynamic learning.  Several physics-informed learning architectures aim at learning hidden dynamics by incorporating physics biases. Deep Lagrangian Networks (DeLaN) [18], for example, injects the Lagrangian mechanics bias into the supervised learning process to improve the performance on difficult trajectories. Sympletic-ODE Net [29], on the other hand, learns the unknown dynamics of a controlled system with Hamiltonian bias. Similar to our framework, Sympletic-ODE exploits the Hamiltonian dynamics for more accurate learning:
@@
Neural ODE.  Similar to our work, Sympletic-ODE Net builds on the Neural ODE architecture [6], which proposes to model a neural network by an ordinary differential equation and learns the rate of change $\frac{dy}{dx}=f(x,y,\theta)$ rather than the function $f$ itself. The backpropagation of NeuralODE is based on the adjoint method with a backward ordinary differential equation on the adjoint states $a(t)=\frac{dL}{dh(t)}$. Comparing with traditional DNNs, Neural ODE [8] allows continuous time-series modeling and doesn’t require discretizing observation and emission intervals.
@@ 1.2 Contributions
We propose novel learning-based frameworks to solve general optimal control problems. Our frameworks actively incorporates Pontryagin maximum principle (PMP) into the learning process. Our contributions include:
@@
    A two-phase-learning framework called NeuralPMP (see section 4) that actively uses Pontryagin maximum principle in the training loss function and the learning process. The two phases utilize a variational autoencoder to prevent static learning while enabling the model to learn the correct Hamiltonian and the corresponding dynamics.
@@ 1.3 Organization
The rest of the paper is organized as follows. Section $2$ briefly introduces the mathematical backgrounds including: optimal control formulations, the Pontryagin maximum principle (PMP), and the definitions of reduced Hamiltonian. Section $3$ proposes an algorithm incorporating PMP into the training loss to solve discrete-time control problems. Sections 4 builds on the limitation of such approach in Section $3$ and introduces a two-phase-learning framework called NeuralPMP for continuous-time optimal control. Section $5$ provides experimental results of NeuralPMP on classical control tasks. Finally, section $6$ concludes our work and outlines future research directions.
@@ 2.1 Optimal control formulation
For simplicity, we work on Euclidean spaces. Our concept can be generalized to a general Banach space. Assume we have state space $X$, a bounded region of $\mathbb{R}^{m}$, with dual $X^{*}$, which can be identified with $\mathbb{R}^{m}$, and a control space $U$, which is another bounded region of $\mathbb{R}^{n}$. Here $m\in\mathbb{N}$ and $n\in\mathbb{N}$ are the state space and control space dimensions. Given a terminal time $T$, we need to find an optimal control path (also called policy) $u^{*}$ so that $u^{*}(t)\in U$ for all time $t\in[0,T]$ and such $u^{*}$ minimizes the following cost functional:
@@ 2.3 Pontryagin maximum principle
The Pontryagin maximum principle [14] provides the necessary condition for an optimal solution of the general optimal control problem.
Theorem 2.1.
@@ 3 Discrete-time Hamiltonian dynamics learning
Before jumping into our main framework in section 4, we propose a discrete-time framework to learn the discretized reduced Hamiltonian dynamics. Moreover, we analyze its advantage over the popular policy-based methods on a specific, irregular linear quadratic control problem.
@@ 3.1 Discrete framework with PMP-based loss
Problem. Suppose an agent $\mathcal{A}$, initially starts at $q_{0}$, needs to perform a sequence of actions $u_{0},u_{1},\cdots,u_{T-1}$ at times $t_{0}=0,t_{1},\cdots,t_{T-1}$ in order to maximize the cost functional $J(q,u)$ introduced in eq. 6.
Solution. From the theory in section 2, we build a parameterized neural network $F_{\theta}$ with parameter $\theta$ that maps from initial state $q_{0}$ to the sequence of optimal actions $(u_{0},u_{1},u_{2},\cdots,u_{T-1})$ by minimizing a Pontryagin maximum principle-based loss function. The training includes 3 steps:
@@
    The adjoint variables $p_{k}$ are then calculated backwards in time $p_{k-1}=p_{k}+\Delta_{k}H_{q}(q_{k},u_{k},p_{k})$, where $H$ is the Hamiltonian defined in eq. 8.
@@
The dynamical system equations eq. 10 and eq. 11 are effectively discretized in the last two steps of the forward pass. We train a network that preserves the last condition eq. 12 of Pontryagin maximum principle (PMP): $\partial_{u}H(x(t),p(t),u(t))=0$. As a result, the loss function for the (policy) network $F_{\theta}$ is defined as:
$L(\theta)=\sum_{k=0}^{T}\lVert H_{u}(q_{k},u_{k},p_{k})\rVert$ (16)
@@ 3.2 Special linear quadratic control problem
subject to the linear dynamics $\dot{a}=Aa+Bu$. For the special case where state space $X=\mathbb{R},A=0,B=1,R=1,Q=0$, the LQR problem is equivalent to the task when a learning agent needs to control its velocity to get as close to the origin as possible even though its initial position can be far from 00. We train the policy $F_{\theta}$ based on the PMP loss on $100$ iterations with 1600 random samples per iteration. The discrete-time horizon is chosen to be $[0,1,2,4,8]$. Notice that the times are not equally spaced, and the values of timestamps are taken into the account in this model. The plot (see fig. 1a) shows that the training process converges after the $70$th iteration.
With the actions produced from the trained policy $F_{\theta}$, the trained car agent quickly decreases the velocity in order to get to 00 as quickly as possible at time 00. At the subsequent times, the car learns to stay where it was. This is the optimal behavior given that the car needs to minimize its cumulative distance to 00 over the time period from 00 to $8$ despite its starting position (see fig. 1b).
Policy-based algorithms in reinforcement learning learn a policy function $\pi$ mapping from $q_{t}$ to the optimal action $u_{t}$. In this special case, we show that no such reasonable function $\pi$ exists. Assume by contradiction, we have such a function $\pi$ so that $u_{t}=\pi(q_{t})$ for each $t$. For $t\geq 1$, $\pi(q_{t})=u_{t}$ must be 00 because once the car gets to the origin, it should stay there forever. This implies that the policy function is approximately zero function. However, the car must increase or decrease its speed to get to 00 as quickly as possible right after it starts. Thus, for a starting value $q_{0}\neq 0$, $u_{0}$, which is approximated by $\pi(q_{0})=0$, needs to be non-zero, leading to a contradiction.
@@ 4 Continuous Hamiltonian dynamics learning with forward-backward variational autoencoder
The training framework in section 3 allows us to gain important insights into the process of actively incorporating PMP principles into the deep learning framework. Nonetheless, the mapping $F_{\theta}$ introduced earlier can be very costly for the time horizon with many steps $t_{1},\cdots,t_{T}$ for large $T$, as it maps to the entire sequence of actions. Moreover, even though the framework can handle uneven time steps, it still requires the discretization of the underlying dynamics, and may not be applicable to a general continuous-time controlled dynamical system. In this section, by also incorporating PMP conditions into the training process, we build a learning framework for continuous-time optimal control problem with a focus on learning the reduced Hamiltonian function eq. 13 and the corresponding Hamiltonian flow.
@@ 4.1 Reduced Hamiltonian learning with PMP-based training loss
    A $P$-net $P_{\phi}$ that takes input as the starting state $q_{0}$ and outputs the adjoint variable $p_{0}$. This $p_{0}$ is necessary since the Hamiltonian $h_{\theta}$ also needs an adjoint variable as input in addition to the state variable.
@@
Training: The following training procedure is called $\textit{Forward}(T)$ for a fixed terminal time $T$. We first create a replay memory $\mathcal{M}$ that consists of tuples with 5 elements $(q,p,\hat{u},\hat{f}(q,\hat{u}),l(q,\hat{u}))$. We then train the reduced Hamiltonian net $h_{\theta}$ and the $P$-net $P_{\phi}$ by repeating the following steps:
@@
    Use NeuralODE [6] to sample the generalized coordinates $(q_{t},p_{t})_{t\geq 0}$ from the Hamiltonian flow defined by equations eq. 14 and eq. 15 with $h=h_{\theta}$. Then for each generalized coordinates $(q_{t},p_{t})$ on the Hamiltonian flow, the 5-tuple $(q_{t},p_{t},\hat{u}(q_{t},p_{t}),\hat{f}(q_{t},\hat{u}_{t}),l(q_{t},\hat{u}_{t}))$ is added to $\mathcal{M}$.
@@
    where $p_{T}$ is the terminal adjoint variable after running $(q,P_{\phi}(q))$ through Hamiltonian dynamics specified by $h_{\theta}$. The variables $\alpha_{1},\alpha_{2}$ and $\beta_{1}$ are hyper-parameters and are optimizable by several techniques: Bayesian optimization [24], Tree-Structured Parzen estimator [3], or bandit algorithms [16]. For efficiency and simplicity, we currently employ grid search with early termination. The first two terms are based on the PMP condition eq. 9, while the last term is used to learn the reduced Hamiltonian correctly.